Hayhayz (hayhayz):

A particular planet has a moment of inertia of 9.74 × 1037 kg•m2 and a mass of 5.98 × 1024 kg. Based on these values, what is the planet’s radius?

Elsa213 (elsa213):

No answer choices? e.e

Hayhayz (hayhayz):

nope ;-; gotta show work

Elsa213 (elsa213):

ew ;-; Give me a seconddd

OpenStudy (goodfaze123):

i'm stumped

OpenStudy (goodfaze123):

i think i got an answer

Hayhayz (hayhayz):

Ummm I think if you times 9.74 by 10^37 you get 9.74 for I but idk what you put after that number

OpenStudy (goodfaze123):

(9.74 x 10^37 kg-m²) = (2/5)(5.98 x 10^24 kg)R² R = 6.38 x 10^6 m

Hayhayz (hayhayz):

Thank you I might need help with one more

OpenStudy (goodfaze123):

did it work out

Hayhayz (hayhayz):

It wont get graded until tomorrow :p

OpenStudy (goodfaze123):

oh

OpenStudy (goodfaze123):

wat is your pic all about it seems really interesting

Hayhayz (hayhayz):

Ah I just found it from a song I listen to, its the cover picture I guess :p

OpenStudy (goodfaze123):

what is the song

Hayhayz (hayhayz):

Nightcall - Dead V from the channel MrSuicideSheep

OpenStudy (goodfaze123):

cool i will check it out

Elsa213 (elsa213):

I created the song...

Hayhayz (hayhayz):

Hey i have 1 more question i will tag you guys

OpenStudy (goodfaze123):

no way